# Opinion Dynamics for Utility Maximizing Agents: Exploring the Impact of Resource Penalty

### v1

@@ 3 Modeling and Problem Setup
$S_{i}(z_{i},p_{i})$ $:=-\frac{w_{i}r_{i}}{B}\left[z_{i}-p_{i}\right]-\frac{z_{i}^{3}}{r_{i}},$ (3b)
$C_{i}(\mathbf{z})$ $:=\sum_{k\in\mathcal{V}}\frac{a_{ik}r_{k}}{B}[z_{k}-z_{i}].$ (3c)
Note that $\forall i\in\mathcal{V}$, the self function $S_{i}(\cdot,\cdot)$ depends only on $i$’s own opinion, its preference $p_{i}$ and other parameters. On the other hand, $\forall i\in\mathcal{V}$, the crowd function $C_{i}(\cdot)$ depends on the deviations of $i$’s opinion $z_{i}$ from its neighbors’ opinions $z_{k}$’s. If $a_{ik}>0\>(<0)$ then $i\in\mathcal{V}$ wants to agree (disgree) with agent $k$ and hence the term in $C_{i}(\cdot)$ corresponding to agent $k$ drives $z_{i}$ towards (away from) the opinion of $k\in\mathcal{V}$.
Now, we make some observations about the self function $S_{i}(\cdot,p_{i})$, which hold $\forall i\in\mathcal{V}$. $S_{i}(\cdot,p_{i})$ is continuous and strictly decreasing with $\lim_{z_{i}\rightarrow-\infty}S_{i}(z_{i},p_{i})=\infty$ and $\lim_{z_{i}\rightarrow+\infty}S_{i}(z_{i},p_{i})=-\infty$. Thus, $S_{i}(\cdot,p_{i})$ has exactly one real root for every fixed value of $p_{i}\in\mathbb{R}$. Let us denote the real root of $S_{i}(\cdot,p_{i})$ as $m_{i}(p_{i})\in\mathbb{R}$, i.e., $S_{i}(m_{i}(p_{i}),p_{i})=0$. Hereafter, we will exclude the preference argument in $S_{i}(\cdot,p_{i})$ and $m_{i}(\cdot)$ wherever there is no confusion. Moreover, by considering $S_{i}(0)$ and $S_{i}(p_{i})$, we can verify that $0\leq|m_{i}|\leq|p_{i}|$ and $m_{i}p_{i}\geq 0$. $m_{i}\in\mathbb{R}$ can be interpreted as the opinion that agent $i$ would attain under (2) if it were socially closed from the influence of other agents, i.e. $a_{ik}=0$, $\forall k\in\mathcal{V}$.
@@
    (No antagonistic relations). For each agent $i\in\mathcal{V}$, $\mathcal{N}_{i}^{e}=\varnothing$. $\bullet$
@@
where
$C_{i}^{+}(\mathbf{z})$ =$=\begin{cases}0\>&;\>\>\text{if}\>\>\mathcal{N}_{i}=\varnothing,\\ \displaystyle{\frac{\sum_{j\in\mathcal{V}}a_{ij}r_{j}}{B}}\,\big{[}\bar{z}_{i}-z_{i}\big{]}\>&;\>\>\text{otherwise}\end{cases}$ (5)
$\bar{z}_{i}$ $:=\sum_{k\in\mathcal{V}}\frac{a_{ik}r_{k}}{\left(\sum_{j\in\mathcal{V}}a_{ij}r_{j}\right)}z_{k}\,.$ (6)
@@
The overall objective of this work is to capture the effect of limited resources available to the agents on the evolution of opinions. The important analytical questions regarding the opinion dynamics (2) that we study in this paper include
@@ 4 Asymptotic Behavior of Opinions
$\mathcal{E}(\mathbf{p}):=\{\mathbf{z}\in\mathbb{R}^{n}\,|\,\dot{\mathbf{z}}=\mathbf{f}(\mathbf{z},\mathbf{p})=\mathbf{0}\}.$ (8)
@@
(Ultimate boundedness of opinions). Let $\mathbf{z}(t)$ be the solution to (2) from the initial condition $\mathbf{z}(0)$. Then $\exists\,\,\eta\geq 0$ (independent of $\mathbf{z}(0)$) and $\exists\,\,T(\mathbf{z}(0))\geq 0$ such that $|z_{i}(t)|\leq\eta$, $\forall i\in\mathcal{V}$, $\forall t\geq T$.
Proof 4.2.
Consider the following radially unbounded positive definite Lyapunov-like function $V:=V(\mathbf{z})=0.5\,\mathbf{z}^{\top}\mathbf{z}$. The time derivative of $V$ along the trajectories of (2) is
$\dot{V}$ $=\sum_{i\in\mathcal{V}}z_{i}\Big{[}-\frac{w_{i}r_{i}}{B}[z_{i}-p_{i}]\Big{]}-\sum_{i\in\mathcal{V}}z_{i}\bigg{[}\sum_{k\in\mathcal{V}}\frac{a_{ik}r_{k}}{B}\left[z_{i}-z_{k}\right]\bigg{]}$
$\quad-\sum_{i\in\mathcal{V}}\frac{z_{i}^{4}}{r_{i}}\leq\sum_{i\in\mathcal{V}}\left[\alpha_{i}|z_{i}|+\beta_{i}|z_{i}|^{2}-\gamma_{i}|z_{i}|^{4}\right]\,$ (9)
for some positive constants $\{\alpha_{i},\beta_{i},\gamma_{i}\}_{i\in\mathcal{V}}$. As the fourth-degree terms dominate the linear and quadratic terms for large enough $\mathbf{z}$, we see that $\exists\zeta\geq 0$ such that $\dot{V}<0$, $\forall\mathbf{z}\in\mathbb{R}^{n}\setminus\Omega_{\zeta}$, where $\Omega_{\zeta}:=\{z\in\mathbb{R}^{n}\>\mid V(\mathbf{z})\leq\zeta\}$. Hence, the set $\Omega_{\zeta}$ is positively invariant under the dynamics (2) and the solution with any initial condition $\mathbf{z}(0)\in\mathbb{R}^{n}$ enters the set $\Omega_{\zeta}$ in finite time $\bar{T}(\mathbf{z}(0))$. Positive-definiteness and radial unboundedness of $V$ then implies the result.
Note that one can explicitly find an ultimate bound $\eta$ but we skip it for brevity. Also note that Theorem 4.1 holds for all values of the parameters and hence does not depend on the type of interactions among agents or the structure of the social network. From the proof of Theorem 4.1, also note that for any initial condition, the solution is uniformly bounded over all time. Then the local Lipschitzness of the vector field in (2) can be used to show existence and uniqueness of solutions of (2) for all time. Next, we characterize the conditions under which the opinions converge to an equilibrium. Recall that for each $i\in\mathcal{V}$, $S_{i}(\cdot)$ has a unique root at $m_{i}$, i.e. $S_{i}(m_{i})=0$. Using this, we show that the opinion of a socially closed agent (if one exists) converges to its corresponding $m_{i}$.
Lemma 4.3.
(Convergence of opinions of socially closed agents). Consider the dynamics given by (3a). Let $\mathbf{z}(t)$ denote the solution of (2) from an initial condition $\mathbf{z}(0)$. Then, $\lim_{t\to\infty}z_{i}(t)=m_{i}\>;\>\forall i\in\mathcal{V}_{cl}$, where $\mathcal{V}_{cl}$ is the set of socially closed agents, i.e.,
$\mathcal{V}_{cl}:=\{i\in\mathcal{V}\>\mid\>\mathcal{N}_{i}=\varnothing\}\subseteq\mathcal{V}\,.$ (10)
Proof 4.4.
If $\mathcal{V}_{cl}=\varnothing$, then there is nothing to prove. So let $\mathcal{V}_{cl}\neq\varnothing$ and consider an arbitrary but fixed $i\in\mathcal{V}_{cl}$. Then from (3) and the definition of $\mathcal{V}_{cl}$, it is clear that $\dot{z}_{i}=S_{i}(z_{i})$ is decoupled from the opinions of other agents. From (7), it is clear that this scalar differential equation has a globally asymptotically stable equilibrium point at $m_{i}$.
@@
    (Weak antagonistic relations). For each agent $i\in\mathcal{V}$,
    $\frac{w_{i}r_{i}}{B}>\sum_{k\in\mathcal{N}_{i}^{e}}\frac{2|a_{ik}|r_{k}}{B},$
    where recall that $\mathcal{N}_{i}^{e}$ is the set of enemies of agent $i\in\mathcal{V}$. $\bullet$
Notice that under Assumption (SA1), $w_{i}>0$ and $r_{i}>0$ for each agent $i\in\mathcal{V}$. Thus, the case of no antagonistic relations stated in Assumption (A3) is a special case of Assumption (A4). All the results which hold under Assumption (A4) also hold under Assumption (A3). Now, we present the main result of this section. Its proof is in the appendix.
Theorem 4.5.
(Existence and uniqueness of equilibrium points). Consider the dynamics given by (2). Suppose Assumption (A4) holds. Then, there exists a unique globally exponentially stable equilibrium point $\mathbf{z}^{*}\in\mathcal{E}$. normal-$\bullet$
Remark 4.6.
(On the weak antagonistic relationships condition). The condition of Assumption (A4) means that the willingness of every agent $i\in\mathcal{V}$ to hold an opinion close to its internal preference is greater than twice the aggregate influence weight of its enemies. Under this assumption, Theorem 4.5 shows that the opinions converge to a unique equilibrium. normal-$\bullet$
@@
Proposition 4.7.
(Convergence to the set $\mathcal{M}^{n}$ in absence of antagonistic relations). Consider the opinion dynamics given by (2). Suppose Assumption (A3) holds. Let $m_{\min}$, $m_{\max}$ and $\mathcal{M}$ be as defined in (11) and (12) respectively. Then, $\mathcal{M}^{n}$ is positively invariant under the opinion dynamics (2). Let $\mathbf{z}(t)$ be the solution to (2) from an initial condition $\mathbf{z}(0)\in\mathbb{R}^{n}$. Then $\mathbf{z}(t)$ converges to $\mathcal{M}^{n}$. Further, define $\mathcal{V}_{\max}:=\{i\in\mathcal{V}\mid m_{i}=m_{\max}\}$ and $\mathcal{V}_{\min}:=\{i\in\mathcal{V}\mid m_{i}=m_{\min}\}$. Suppose $m_{\min}<m_{\max}$. Then the following statements are equivalent.
@@
    $\forall i\in\mathcal{V}_{\max}$, $\exists$ a directed walk in $\mathsf{G}$ starting from $j\in\mathcal{V}\setminus\mathcal{V}_{\max}$ to $i$ and $\forall i\in\mathcal{V}_{\min}$, $\exists$ a directed walk in $\mathsf{G}$ starting from $j\in\mathcal{V}\setminus\mathcal{V}_{\min}$ to $i$. normal-$\bullet$
@@ 5 Consensus and Nash Equilibria
In this section, we analyze consensus equilibrium points of the opinion dynamics (2) and Nash equilibria of the underlying game. We also explore the relation between the Nash equilibrium set of the underlying game and the set of equilibria of the opinion dynamics (2). Finally, we also analyze the price of anarchy.
@@ 5.1 Consensus Equilibria
First, we deal with the consensus equilibria of the model, i.e., equilibria of the form $\xi\mathbf{1}$, with $\xi\in\mathbb{R}$. We refer to the case of $\xi=0$ as a neutral consensus since all the agents have neutral opinions (equal to 00) in this case. On the other hand, we refer to the case of $\xi\neq 0$ as a non-neutral consensus. In the following lemma, we present conditions for (2) to have a consensus equilibrium. We use the form of the dynamics in (3a) and the functions in (3) to justify our claims.
Theorem 5.1.
(Necessary and sufficient conditions for existence of a consensus equilibrium). Consider the dynamics (2) (equivalently (3a)). For each $i\in\mathcal{V}$, let $m_{i}\in\mathbb{R}$ be the unique point such that $S_{i}(m_{i})=0$. Then, $\xi\mathbf{1}\in\mathcal{E}$ if and only if $m_{i}=\xi$, $\forall i\in\mathcal{V}$.
@@
If $z_{i}=\xi$, $\forall i\in\mathcal{V}$, for some $\xi\in\mathbb{R}$ then $C_{i}(\xi\mathbf{1})=0$, $\forall i\in\mathcal{V}$. Hence, from (3a), $\xi\mathbf{1}\in\mathcal{E}$ iff $S_{i}(\xi)=0,\forall i\in\mathcal{V}$. Since $m_{i}$ is the unique root of $S_{i}(.)$, the claim follows.
@@
(Consensus formation among agents). Theorem 5.1 states that it is both necessary and sufficient for all the $m_{i}$’s to be the same for the opinion dynamics model to have a consensus equilibrium. It is evident that if the agents are to arrive at a consensus equilibrium, then all their preferences $p_{i}$’s must be of the same sign. When $p_{i}=0$, $\forall i\in\mathcal{V}$, then the only possible consensus equilibrium is the neutral consensus, i.e., every agent reaches a neutral opinion on the topic. If the preferences of the agents have different signs, then the opinions of agents can never reach an exact consensus in equilibrium. However, other equilibria that are arbitrarily close to consensus may still exist. Now, if the weak antagonistic relationships condition given in Assumption (A4) holds and there exists a consensus equilibrium, then, from Theorem 4.5, it is the only equilibrium of the dynamics and the agents always achieve consensus starting from any initial opinion vector. normal-$\bullet$
When the agents attain a consensus equilibrium, we can measure how much influence an agent has on the whole group by measuring the deviation of the consensus value from its preference. More specifically, if $\xi\in\mathbb{R}$ is the consensus value at equilibrium, we compute for each $i\in\mathcal{V}$,
$\Delta_{i}(\xi):=|p_{i}-\xi|,$ (13)
as a measure of its influence. Note that if $p_{i}=0$, for some $i\in\mathcal{V}$, then $m_{i}=0$ and hence the only consensus equilibrium possible (if it exists) is neutral. So we consider $p_{i}\neq 0$, $\forall i\in\mathcal{V}$ to give the next result on dominance and discuss it in the remark following it.
@@
(Consensus deviation from preference). Consider the dynamics (2) or equivalently (3a). Suppose that $p_{i}\neq 0$, $\forall i\in\mathcal{V}$. For each agent $i\in\mathcal{V}$, let us define $\sigma_{i}:=w_{i}r_{i}^{2}$ and let $\Delta_{i}(\xi)$ be as in (13). Suppose $\xi\mathbf{1}\in\mathcal{E}$, with $\xi\in\mathbb{R}$. Then $\sigma_{i}\Delta_{i}(\xi)=\sigma_{j}\Delta_{j}(\xi)$, $\forall i,j\in\mathcal{V}$. In particular
$\sigma_{i}>\sigma_{j}\iff\Delta_{i}(\xi)<\Delta_{j}(\xi).$
@@
(Dominance in consensus). Let us call $\forall i\in\mathcal{V}$, the scalar $\sigma_{i}$ as the consensus dominance weight of the agent $i$. Suppose all agents have a non-neutral preference and they attain consensus at $\xi\in\mathbb{R}$. Then, Proposition 5.4 states that if an agent $i\in\mathcal{V}$ has higher consensus dominance weight than agent $j\in\mathcal{V}$, then $\xi$ is closer to $i$’s preference than that of $j$. Note that the consensus dominance weight is directly proportional to the importance weight an agent assigns to its preference and the square of the resources available to it. Thus, an agent with very high resources can exert more influence even if it gives less importance weight to its internal preference. On the other hand, if an agent has lower resources, then it has to have much higher importance weight to have more influence in the group. normal-$\bullet$

### v2

@@ 3 Modeling and Problem Setup
$S_{i}(z_{i},p_{i})$ $:=-w_{i}\left[z_{i}-p_{i}\right]-\frac{z_{i}^{3}}{r_{i}},$ (3b)
$C_{i}(\mathbf{z})$ $:=\sum_{k\in\mathcal{V}}a_{ik}[z_{k}-z_{i}].$ (3c)
Note that $\forall i\in\mathcal{V}$, the self function $S_{i}(\cdot,p_{i})$ depends only on $i$’s own opinion, its preference $p_{i}$ and other parameters. On the other hand, $\forall i\in\mathcal{V}$, the crowd function $C_{i}(\cdot)$ depends on the deviations of $i$’s opinion $z_{i}$ from its neighbors’ opinions $z_{k}$’s. If $a_{ik}>0\>(<0)$ then $i\in\mathcal{V}$ wants to agree (disagree) with agent $k$ and hence the term in $C_{i}(\cdot)$ corresponding to agent $k$ drives $z_{i}$ towards (away from) the opinion of $k\in\mathcal{V}$.
Now, we make some observations about the self function $S_{i}(\cdot,p_{i})$, which hold $\forall i\in\mathcal{V}$. $S_{i}(\cdot,p_{i})$ is continuous and strictly decreasing with $\lim_{z_{i}\rightarrow-\infty}S_{i}(z_{i},p_{i})=\infty$ and $\lim_{z_{i}\rightarrow+\infty}S_{i}(z_{i},p_{i})=-\infty$. Thus, $S_{i}(\cdot,p_{i})$ has exactly one real root for every fixed value of $p_{i}\in$. Let us denote the real root of $S_{i}(\cdot,p_{i})$ as $m_{i}(p_{i})\in$, i.e., $S_{i}(m_{i}(p_{i}),p_{i})=0$. Hereafter, we will exclude the preference argument in $S_{i}(\cdot,p_{i})$ and $m_{i}(\cdot)$ wherever there is no confusion. Moreover, by considering $S_{i}(0)$ and $S_{i}(p_{i})$, we can verify that $0\leq|m_{i}|\leq|p_{i}|$ and $m_{i}p_{i}\geq 0$. $m_{i}\in$ can be interpreted as the opinion that agent $i$ would attain under (2) if it were socially closed from the influence of other agents, i.e. $a_{ik}=0$, $\forall k\in\mathcal{V}$.
@@
    (No antagonistic relations.) $\forall i\in\mathcal{V}$, $\mathcal{N}_{i}^{e}=\varnothing$. $\bullet$
@@
where $\text{where }\ C_{i}^{+}(\mathbf{z})$ =$=\begin{cases}0\>&;\>\>\text{if}\>\>\mathcal{N}_{i}=\varnothing,\\ \displaystyle{\sum_{j\in\mathcal{V}}a_{ij}}\,\big{[}\bar{z}_{i}-z_{i}\big{]}\>&;\>\>\text{otherwise}\end{cases}$ (5)
$\bar{z}_{i}$ $:=\sum_{k\in\mathcal{V}}\frac{a_{ik}}{\left(\sum_{j\in\mathcal{V}}a_{ij}\right)}z_{k}\,.$ (6)
@@
The important analytical questions regarding the opinion dynamics (2) that we study in this paper include
@@ 4 Asymptotic Behavior of Opinions
$\mathcal{E}(\mathbf{p}):=\{\mathbf{z}\in^{n}\,|\,\dot{\mathbf{z}}=\mathbf{f}(\mathbf{z},\mathbf{p})=\mathbf{0}\}.$ (8)
@@
(Ultimate boundedness of opinions.) Let $\mathbf{z}(t)$ be the solution to (2) from the initial condition $\mathbf{z}(0)$. Then $\exists\,\,\eta\geq 0$ (independent of $\mathbf{z}(0)$) and $\exists\,\,T(\mathbf{z}(0))\geq 0$ such that $|z_{i}(t)|\leq\eta$, $\forall i\in\mathcal{V}$, $\forall t\geq T$. Additionally, $\exists\>\Omega\subseteq\mathbb{R}^{n}$ which is convex, compact and positive invariant under (2). $\bullet$
The existence of an ultimate bound can be shown using similar arguments involved in the proof of [31, Theorem 3.1]. Using the same arguments, we can also show existence of a convex, compact sublevel set $\Omega$ of the quadratic Lyapunov function $V(\mathbf{z}):=0.5\;\mathbf{z}^{\top}\mathbf{z}$ such that $\dot{V}(\mathbf{z})<0$ outside $\Omega$ and the solution with any initial condition $\mathbf{z}(0)\in^{n}$ enters the set $\Omega$ in finite time $\bar{T}(\mathbf{z}(0))$. We skip the proof for brevity.
Note that one can explicitly find an ultimate bound $\eta$ but we skip it for brevity. Also note that Theorem 4.1 holds for all values of the parameters and hence does not depend on the type of interactions among agents or the structure of the social network. From the proof of Theorem 4.1, we note that for any initial condition, the solution is uniformly bounded over all time. Then the local Lipschitzness of the vector field in (2) can be used to show existence and uniqueness of solutions of (2) for all time. Next, we characterize the conditions under which the opinions converge to an equilibrium. Recall that for each $i\in\mathcal{V}$, $S_{i}(\cdot)$ has a unique root at $m_{i}$, i.e. $S_{i}(m_{i})=0$. Using this, we show that the opinion of a socially closed agent (if one exists) converges to its corresponding $m_{i}$. The next result can be shown using (7) and the fact that the dynamics (2) for every such agent reduces to the scalar differential equation $\dot{z_{i}}=S(z_{i})$. We skip the proof for brevity.
Lemma 4.2
(Convergence of opinions of socially closed agents.) Consider the dynamics given by (3a). Let $\mathbf{z}(t)$ denote the solution of (2) from an initial condition $\mathbf{z}(0)$. Then, $\lim_{t\to\infty}z_{i}(t)=m_{i}\>;\>\forall i\in\mathcal{V}_{cl}$, where $\mathcal{V}_{cl}$ is the set of socially closed agents, i.e., $\mathcal{V}_{cl}:=\{i\in\mathcal{V}\>\mid\>\mathcal{N}_{i}=\varnothing\}\subseteq\mathcal{V}$. $\bullet$
@@
    (Weak antagonistic relations.) For each agent $i\in\mathcal{V}$, $w_{i}>\sum_{k\in\mathcal{N}_{i}^{e}}2|a_{ik}|$, where $\mathcal{N}_{i}^{e}$ is the set of enemies of $i$. $\bullet$
Notice that under Assumption (SA1), $w_{i}>0,\forall i\in\mathcal{V}$. Thus, the case of no antagonistic relations stated in Assumption (A3) is a special case of Assumption (A4). All the results which hold under Assumption (A4) also hold under Assumption (A3). Now, we present the main result of this section. Its proof is in the appendix.
Theorem 4.3
(Existence and uniqueness of equilibrium points.) Consider the dynamics given by (2). Suppose Assumption (A4) holds. Then, there exists a unique globally exponentially stable equilibrium point $\mathbf{z}^{*}\in\mathcal{E}$. $\bullet$
Remark 4.4
(On the weak antagonistic relationships condition.) The condition of Assumption (A4) means that the willingness of every agent $i\in\mathcal{V}$ to hold an opinion close to its internal preference is greater than twice the aggregate influence weight of its enemies. Under this assumption, Theorem 4.3 shows that the opinions converge to a unique equilibrium. $\bullet$
@@
Proposition 4.5
(Convergence to the set $\mathcal{M}^{n}$ in absence of antagonistic relations.) Consider the opinion dynamics given by (2). Suppose Assumption (A3) holds. Let $m_{\min}$, $m_{\max}$ and $\mathcal{M}$ be as defined in (9) and (10) respectively. Then, $\mathcal{M}^{n}$ is positively invariant under the opinion dynamics (2). Let $\mathbf{z}(t)$ be the solution to (2) from an initial condition $\mathbf{z}(0)\in^{n}$. Then $\mathbf{z}(t)$ converges to $\mathcal{M}^{n}$. Further, define $\mathcal{V}_{\max}:=\{i\in\mathcal{V}\mid m_{i}=m_{\max}\}$ and $\mathcal{V}_{\min}:=\{i\in\mathcal{V}\mid m_{i}=m_{\min}\}$. Suppose $m_{\min}<m_{\max}$. Then the following statements are equivalent.
@@
    $\forall i\in\mathcal{V}_{\max}$, $\exists$ a directed walk in $\mathsf{G}$ starting from $j\in\mathcal{V}\setminus\mathcal{V}_{\max}$ to $i$ and $\forall i\in\mathcal{V}_{\min}$, $\exists$ a directed walk in $\mathsf{G}$ starting from $j\in\mathcal{V}\setminus\mathcal{V}_{\min}$ to $i$$\bullet$
@@ 5 Consensus and Nash Equilibria
In this section, we analyze consensus equilibria of the opinion dynamics (2) and Nash equilibria of the underlying game. We also explore the relation between the Nash equilibrium set of the underlying game and the set of equilibria of the opinion dynamics (2). Finally, we also analyze the price of anarchy.
@@ 5.1 Consensus Equilibria
First, we deal with the consensus equilibria of the model, i.e., equilibria of the form $\xi\mathbf{1}$, with $\xi\in$. We refer to the case of $\xi=0$ as a neutral consensus since all the agents have neutral opinions (equal to 00) in this case. On the other hand, we refer to the case of $\xi\neq 0$ as a non-neutral consensus. In the following lemma, we present conditions for (2) to have a consensus equilibrium. We use the form of the dynamics in (3a) and the functions in (3) to justify our claims.
Theorem 5.1
(Necessary and sufficient conditions for existence of a consensus equilibrium.) Consider the dynamics (2) (equivalently (3a)). For each $i\in\mathcal{V}$, let $m_{i}\in$ be the unique point such that $S_{i}(m_{i})=0$. Then, $\xi\mathbf{1}\in\mathcal{E}$ if and only if $m_{i}=\xi$, $\forall i\in\mathcal{V}$.
@@
If $z_{i}=\xi$, $\forall i\in\mathcal{V}$, for some $\xi\in$ then $C_{i}(\xi\mathbf{1})=0$, $\forall i\in\mathcal{V}$. Hence, from (3a), $\xi\mathbf{1}\in\mathcal{E}$ iff $S_{i}(\xi)=0,\forall i\in\mathcal{V}$. Since $m_{i}$ is the unique root of $S_{i}(.)$, the claim follows.
@@
(Consensus formation among agents.) Theorem 5.1 states that it is both necessary and sufficient for all the $m_{i}$’s to be the same for the opinion dynamics model to have a consensus equilibrium. It is evident that if the agents are to arrive at a consensus equilibrium, then all their preferences $p_{i}$’s must be of the same sign. When $p_{i}=0$, $\forall i\in\mathcal{V}$, then the only possible consensus equilibrium is the neutral consensus, i.e., every agent reaches a neutral opinion on the topic. If the preferences of the agents have different signs, then the opinions of agents can never reach an exact consensus in equilibrium. However, other equilibria that are arbitrarily close to consensus may still exist.
Now, if the weak antagonistic relationships condition given in Assumption (A4) holds and there exists a consensus equilibrium, then, from Theorem 4.3, it is the only equilibrium of the dynamics and the agents always achieve consensus starting from any initial opinion vector. $\bullet$
When the agents attain a consensus equilibrium, we can measure how much influence an agent has on the whole group by measuring the deviation of the consensus value from its preference. Note that if $p_{i}=0$, for some $i\in\mathcal{V}$, then $m_{i}=0$ and hence the only consensus equilibrium possible (if it exists) is neutral. So we consider $p_{i}\neq 0$, $\forall i\in\mathcal{V}$ to give the next result on dominance and discuss it in the remark following it.
@@
(Consensus deviation from preference.) Consider the dynamics (2) or equivalently (3a). Suppose that $p_{i}\neq 0$, $\forall i\in\mathcal{V}$. For each agent $i\in\mathcal{V}$, let us define $\sigma_{i}:=w_{i}r_{i}$ and $\Delta_{i}(\xi):=|p_{i}-\xi|$. If $\xi\mathbf{1}\in\mathcal{E}$, with $\xi\in$, then $\sigma_{i}\Delta_{i}(\xi)=\sigma_{j}\Delta_{j}(\xi)$, $\forall i,j\in\mathcal{V}$. In particular, $\sigma_{i}>\sigma_{j}$ iff $\Delta_{i}(\xi)<\Delta_{j}(\xi)$.
@@
(Dominance in consensus.) Let us call $\forall i\in\mathcal{V}$, the scalar $\sigma_{i}$ as the consensus dominance weight of the agent $i$. Suppose all agents have a non-neutral preference and they attain consensus at $\xi\in$. Then, Proposition 5.4 states that if an agent $i\in\mathcal{V}$ has higher consensus dominance weight than agent $j\in\mathcal{V}$, then $\xi$ is closer to $i$’s preference than that of $j$. Note that the consensus dominance weight is directly proportional to the importance weight an agent assigns to its preference and the resources available to it. Thus, an agent with very high resources can exert more influence even if it gives less importance weight to its internal preference. On the other hand, if an agent has lower resources, then it has to have much higher importance weight to have more influence in the group. $\bullet$